### v3

@@ Appendix A Minimax Proof for Minimizing CDL
We can prove the existence of an algorithm that achieves $\tilde{O}(\sqrt{T})$ CDL via the minimax theorem. The minimax theorem demonstrates a remarkable difference between $K_{1}$ and CDL - simply by reporting the adversary’s strategy truthfully, CDL matches the lowerbound.
In this setup, a predictor $F$ makes a prediction $p_{t}\in[0,1]$ at each time step $t=1,2,\ldots$, and an adversary $A$ picks an outcome $\theta_{t}\in\{0,1\}$. Both $p_{t}$ and $\theta_{t}$ are chosen based on the history $h_{t-1}=(p_{1},\theta_{1},p_{2},\theta_{2},\ldots,p_{t-1},\theta_{t-1})$, but these two quantities at the same time step $t$ cannot depend on each other. The goal of the predictor is to minimize $\textsc{CDL}(h_{T})$.
@@
$\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=O(\log(T)\sqrt{T}),$
@@
We prove the following argument which implies Theorem A.1 immediately:
$\min_{{\mathcal{F}}}\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=O(\log(T)\sqrt{T}),$
@@
The following holds:
$\max_{{\mathcal{A}}}\min_{F}\operatorname{\mathbb{E}}_{A\sim{\mathcal{A}}}[\textsc{CDL}(h_{F,A})]\leq O(\log(T)\sqrt{T}).$
@@
${\mathbf{E}}\left[\textsc{CDL}(\bm{p},\bm{\theta})\right]$ $\leq 4m{\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]+4\alpha\sqrt{T}+4\beta T+O(\alpha^{2}m\log m)$
$\leq O(\sqrt{T}/\log(T))+O(\sqrt{T\log(T)})+O(\sqrt{T}\log T)+O(\sqrt{T}\log(T))$
$=O(\sqrt{T}\log T).\qed$

### v4

@@ Appendix A Minimax Proof for Minimizing CDL
We prove the existence of an algorithm that achieves $O(\frac{\log T}{\sqrt{T}})$ CDL via the minimax theorem. The minimax theorem allows us to assume that the adversary’s (randomized) strategy is fixed and known by the predictor. Our proof demonstrates a remarkable difference between ECE and CDL - simply by truthfully reporting the mean of each state (conditioned on the history and rounded to a suitable finite subset of $[0,1]$), our predictor achieves the $O(\frac{\log T}{\sqrt{T}})$ CDL rate that nearly matches the natural $\Omega(\frac{1}{\sqrt{T}})$ lower bound.
In this setup, a predictor $F$ makes a prediction $p_{t}\in[0,1]$ at each time step $t=1,2,\ldots$, and an adversary $A$ picks an outcome $\theta_{t}\in\{0,1\}$. Both $p_{t}$ and $\theta_{t}$ are chosen based on the history $h_{t-1}=(p_{1},\theta_{1},p_{2},\theta_{2},\ldots,p_{t-1},\theta_{t-1})$. The goal of the predictor is to minimize $\textsc{CDL}(h_{T})$.
@@
$\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)},$
@@
We prove the following statement which implies Theorem A.1 immediately: there exists a positive integer $m$ such that
$\min_{{\mathcal{F}}}\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)},$
@@
There exists a positive integer $m$ such that
$\max_{{\mathcal{A}}}\min_{F}\operatorname{\mathbb{E}}_{A\sim{\mathcal{A}}}[\textsc{CDL}(h_{F,A})]\leq O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}.$
@@
$\textsc{CDL}(\bm{p},\bm{\theta})$ $\leq\frac{4m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{4\alpha}{\sqrt{T}}+4\beta+O\Big{(}\frac{\alpha^{2}m\log m}{T}\Big{)}.$
$\leq O\Big{(}\frac{1}{\sqrt{T}\log T}\Big{)}+O\Big{(}\frac{\sqrt{\log T}}{\sqrt{T}}\Big{)}+O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}+O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}$
$=O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}.\qed$